### 21 Dec 2023 v3

# Heterogeneous Position Effects and the Power of Rankings

Abstract
Many online retailers and search intermediaries present products on ranked product lists. Because changing the ordering of products influences their revenues and consumer welfare, these platforms can design rankings to maximize either of these two metrics. In this paper, I study how rankings following these two objectives differ and what determines this difference. First, I highlight that when position effects are heterogeneous, rankings can increase both revenues and consumer welfare by increasing the overall conversion rate. Second, I provide empirical evidence for this heterogeneity, highlighting that cheaper alternatives have stronger position effects. Finally, I quantify revenue and consumer welfare effects across the different rankings. To this end, I develop an estimation procedure for the search and discovery model of Greminger (2022) that yields a smooth likelihood function by construction. The results from the model show that rankings designed to maximize revenues can also increase consumer welfare. Moreover, these revenue-based rankings reduce consumer welfare only to a limited extent relative to a consumer-welfare-maximizing ranking.
Introduction
Many consumers purchase products from online retailers or search intermediaries. These online platforms typically present consumers with ranked product lists that have been shown to entail substantial position effects: products shown in higher positions are more often clicked on and bought (e.g., Ghose et al., 2014; De Los Santos and Koulayev, 2017; Ursu, 2018). These position effects allow online platforms to increase their revenues by changing the ranking (e.g., Choi and Mela, 2019). At the same time, changing the ranking also affects consumers; it can help them find better-matching alternatives and save them time and effort going down the list. Hence, platforms concerned about customer churn can also use rankings to make consumers better off (e.g., Ursu, 2018). However, following one of these objectives potentially comes at the expense of the other. A ranking that maximizes the platform’s revenues could end up harming consumers, whereas a consumer-welfare-maximizing ranking may reduce the platform’s revenues.
In this paper, I ask whether this is the case and study how rankings following the two objectives differ in how they affect consumers and a platform’s revenues from a price-based commission. First, I use a simple example to highlight a crucial factor determining such differences: heterogeneous position effects. If there are position effects, a ranking can shift demand to more expensive alternatives by moving them higher on the list. However, suppose position effects are heterogeneous and cheaper alternatives have stronger position effects. In this case, moving expensive alternatives up on the list also decreases the conversion rate such that maximizing the platform’s revenues requires balancing the two effects. Hence, depending on whether position effects are heterogeneous, a revenue-maximizing ranking will move more or less expensive alternatives to the top and affect consumers accordingly.
Whether cheaper alternatives have stronger position effects such that a platform needs to balance the two effects is an empirical question and not clear a priori. To address it, I use experimental ranking variation to identify heterogeneous position effects in data from a large online search intermediary, Expedia. The results show that cheaper alternatives indeed have stronger position effects. More broadly, I find that product attributes and the position on the list complement each other such that hotels with desirable attributes have stronger position effects. Hence, moving hotels with a lower price and other desirable attributes higher up on the list will increase the conversion rate on average, potentially increasing both revenues and consumer welfare.
Analyzing whether this is the case requires quantifying how the different rankings affect platform revenues and consumer welfare. To this end, I build on and estimate the search and discovery (SD) model of Greminger (2022). In the model, consumers decide between discovering more products by scrolling down the list and learning more about an already-discovered product by going to its detail page. I argue that this model is well-suited to study the consumer welfare effects of different rankings because it provides a micro-foundation for position effects: consumers decide to stop scrolling before ever reaching the lower positions on the list. Moreover, I show that this discovery mechanism implies position effects that closely match the pattern revealed by the descriptive evidence: alternatives with desirable attributes have stronger position effects. As a result, the estimated model predicts heterogeneous position effects that closely match those in the data and, hence, fits the moments determining the trade-off between the different rankings notably well.
Building on the model, I introduce two heuristics for the revenue-maximizing ranking, the “bottom-up” and the “Position-1” ranking, and analyze their effects in a counterfactual analysis. The results show that these revenue-based rankings increase total revenues by up to 14.93% over a randomized ranking. Notably, this revenue increase does not come at a cost to consumers. Instead, the proposed rankings increase consumer welfare over a randomized ranking by up to 0.92%, conditional on consumers searching at least one alternative. These consumer welfare gains come close to the 0.98% gain achieved by the utility-based ranking, which I prove to maximize consumer welfare. These results show that revenue-based rankings are closer to the consumer-welfare-maximizing ranking than to a randomized ranking due to how heterogeneous position effects stem from consumers’ search decisions. Hence, both consumers and the platform can benefit from rankings that maximize revenues.
By studying the effects of different rankings, this paper also offers two managerial insights. First, a concern for online platforms is that revenue-based rankings could lead to a bad search experience for consumers, thus increasing customer churn. My results suggest that this is not the case. Because high-utility alternatives have stronger position effects, revenue-based rankings increase conversions by moving them higher up on the list, thus also benefiting consumers relative to various other rankings. Second, I also show that a utility-based ranking maximizes consumer welfare and the overall purchase probability. Hence, platforms that explicitly want to increase consumer welfare can do so by maximizing conversions, an observable outcome that ranking algorithms can directly target.
To address the substantive research question, I also make several methodological contributions. First, I develop an estimation approach for the SD model that yields a smooth likelihood by construction and avoids computationally expensive steps during estimation. Second, I extend the analysis of Greminger (2022) to show how the SD model captures position effects conditional on the observed product attributes. Finally, I use this result to show how the observed position effects are sufficient to identify model parameters without observing how many products consumers discover.
Related Literature
This paper contributes to the literature examining the effects of different rankings on consumer welfare, search behavior, and platform revenues. Initial studies in this literature focus on the effects of utility-based rankings (Ghose et al., 2012, 2014; Ursu, 2018). More recent studies additionally consider the effects of rankings designed to increase platform revenues, documenting possible trade-offs between making consumers better off and increasing platform revenues (Choi and Mela, 2019; Zhang et al., 2021; Compiani et al., 2023; Donnelly et al., 2023). I add to this literature by highlighting how this trade-off stems from heterogeneous position effects, deriving how these heterogeneous position effects arise in a micro-founded model, and quantifying this trade-off by estimating the same model. By focusing on the effects of rankings with different objectives, this paper also differs from recent work by Zhang et al. (2023), who estimate the SD model to study how different search routes affect consumers’ search behavior.
By providing empirical evidence that shows how position effects depend on product attributes like price, I also add to the empirical literature that documents position effects on ranked product lists (Ghose et al., 2012, 2014; De Los Santos and Koulayev, 2017; Ursu, 2018) and in search advertising (Goldman and Rao, 2014; Athey and Imbens, 2015; Jeziorski and Segal, 2015; Jeziorski and Moorthy, 2018). [^1] These studies find that ads with a larger click-through-rate on any position have weaker position effects. The difference to my results suggests that consumers use a general-purpose search engine like Google with a specific outcome already in mind instead of using the list to compare different alternatives. This is corroborated by results of Blake et al. (2015), who find that search advertising is not effective for a well-known platform like eBay. Unlike Ghose et al. (2014), who also consider how position effects vary with product attributes, I use experimental variation in rankings to account for endogeneity in rankings and consistently find that desirable attributes amplify position effects.
The methodological contributions in this paper add to the empirical search literature that has used a range of alternative models to rationalize position effects (e.g., Chan and Park, 2015; Chen and Yao, 2017; De Los Santos and Koulayev, 2017; Ursu, 2018; Choi and Mela, 2019; Gibbard, 2022b; Compiani et al., 2023; Donnelly et al., 2023). The SD model I estimate has several features that make it particularly useful to capture position effects and quantify the effects of different rankings, and I highlight these features and how they relate to other sequential search models after introducing the model. As they are nested in the SD model, my estimation approach can also be used to estimate directed search models á la Weitzman (1979) and, hence, adds to the recent literature introducing estimation approaches without researcher-chosen smoothing parameters for these models Jiang et al. (2021); Chung et al. (2023). However, unlike these approaches, my approach does not require observing the search order, is easier to implement, and requires fewer computations. [^2] The approaches of Jiang et al. (2021) and Chung et al. (2023) grow in computational complexity in the number of clicks each consumer makes, which is not the case with my approach.
By showing that a utility-based ranking maximizes consumer welfare in the SD model and introducing two heuristics for maximizing revenues in the short-term, I also add to the literature deriving algorithms to optimally choose in which order to display alternatives to consumers (e.g., Ryzin and Mahajan, 1999; Talluri and Van Ryzin, 2004; Davis et al., 2014; Ursu and Dzyabura, 2020; Chu et al., 2020; Derakhshan et al., 2022; Compiani et al., 2023). Whereas this literature motivates this paper’s empirical evaluation of how revenue-maximizing rankings affect consumers, it tends to rely on more stylized demand models and, [^3] Early work in this literature did not consider search. The model of Chu et al. (2020) precludes position effects for consumers who do not buy an alternative, contrasting what I observe in the data. Derakhshan et al. (2022) consider heterogeneous position effects in the context of maximizing the conversion rate (as opposed to revenues) and assume that consumers first decide how many products to discover before inspecting all of the discovered products at the same time. This model also contrasts the data as almost all consumers skip clicking on some alternatives. Ursu and Dzyabura (2020) instead focus on a retailer deciding how to place independent categories within a retailer. except for Compiani et al. (2023), does not quantify the different effects.
Heterogeneous Position Effects and Ranking Objectives
To motivate the empirical analysis, I first use a simple example to highlight a crucial factor determining how a revenue-maximizing ranking orders alternatives and, hence, how such a ranking will affect consumers.
Consider an online retail platform that offers two products, $A$ and $B$. $B$ is more expensive ($p_{B}>p_{A}$), and the platform decides whether to move it to the first position. The platform takes the same price-based commission from both products and, therefore, wants to maximize revenues across both positions. Hence, whether the platform will display $B$ first depends on how total revenues change following the switch. This change is given by
[…]
where $p_{j}$ denotes the price of alternative $j$ and $\Delta d_{j}$ denotes its position effect, defined as the absolute demand change when moving from the second to the first position. In line with ample empirical evidence (e.g. Ursu, 2018), these position effects will be non-negative ($\Delta d_{j}\geq 0\forall j$).
The two parts in equation (1) reveal that changing the ranking can increase revenues by shifting demand to more expensive alternatives through a “price effect”, increasing the conversion rate through a “demand effect”, or a combination of both. Crucially, the demand effect reveals that heterogeneous position effects play an important role in determining which alternatives a revenue-maximizing ranking will move to the top and, hence, how it will affect consumers. If position effects are homogeneous ($\Delta d_{A}=\Delta d_{B}$), there is no demand effect such that moving the more expensive alternative $B$ to the top necessarily increases revenues. In contrast, if the cheaper alternative $A$ has a stronger position effect ($\Delta d_{A}>\Delta d_{B}$), the demand effect will be negative such that choosing a ranking to maximize revenues becomes similar to a monopolist pricing problem.
Which of the two cases applies and how the resulting ranking affects consumers is not clear a priori. For example, suppose the two products are homogeneous and consumers never choose the alternative in the second position because the costs of looking at the second position are prohibitively high. If consumers are not too price-sensitive and choose whichever product is in the first position, then $\Delta d_{B}=\Delta d_{A}>0$ such that moving $B$ to the top maximizes revenues while harming consumers who are now getting the more expensive alternative. Now suppose that consumers instead are so price-sensitive that they will not buy $B$ even when it is shown in the first position. In this case, $\Delta d_{A}>\Delta d_{B}=0$ such that moving $A$ to the top maximizes revenues while also maximizing consumer welfare. Finally, suppose products are not homogeneous and all consumers prefer $B$ to $A$ despite the higher price, while only some consumers are not willing to buy $A$. In this case, $\Delta d_{B}>\Delta d_{A}>0$ such that moving $B$ to the top maximizes revenues and consumer welfare.
As these examples highlight, heterogeneous position effects determine how the different rankings will affect consumers and the platform’s revenues. [^4] For simplicity, I focus only on two products. With more than two products, the revenue change would also depend on demand changes for alternatives other than the ones being moved. Nonetheless, the logic would be largely the same and rankings would continue to increase revenues by shifting demand to more expensive alternatives or increasing conversions, with the latter being mainly determined through heterogeneous position effects. Moreover, they can be informative of consumers’ preferences and how they interact with ranked product lists. Hence, in a first step to show how the different rankings affect consumers and platform revenues, I now provide empirical evidence that shows whether position effects are heterogeneous and how they vary with product attributes like price.
Data and Descriptive Evidence
I use click-stream data from a leading online travel agency, Expedia. The data can be obtained from Kaggle.com and contains information on clicks and purchases for 166,039 search sessions between November 2012 and June 2013.555The data is available under the following link: https://www.kaggle.com/c/expedia-personalized-sort/data. A search session starts at the point where a consumer has submitted a query for a hotel stay on Expedia. Following this query, Expedia presents a list of the available options on which consumers observe various hotel characteristics, such as the price per night and the review score. They then can scroll down to reveal more hotels on the list or click on a particular item they already see on the list. Clicking on an item leads to the hotel’s detail page, which reveals further information and allows booking the hotel. Ursu (2018) provides a more comprehensive discussion of the dataset, and I apply similar criteria to prepare the final sample (see Appendix E).
Summary
The main feature of the data is that for about 30% of search sessions, Expedia randomly assigned hotels that fit the query to positions on the list. This exogenous variation is essential to identify position effects without convoluting the effect of more desirable hotels being ranked higher. Suppose hotels instead are positioned on top of the list based on unobservable characteristics that also lead to more clicks and purchases. In that case, it is challenging to disentangle correlations with such unobservables, potentially leading to an overestimation of position effects (Ursu, 2018). For the other 70% of consumers in the sample, Expedia used its ranking algorithm to assign hotels to positions. [^6] It is not possible to get the specifics of the Expedia ranking algorithm used during the sample period. Most likely, it used observable hotel and query characteristics to create a ranking that maximized predicted clicks or bookings.
Table 1 summarizes the dataset on a hotel and session level for consumers who observed the randomized ranking.  [^7] Appendix E provides a detailed description of each variable. The “no reviews” variable is a dummy indicating whether a hotel has no reviews. This is coded as a “review score” of zero in the raw data. However, given that it differs from a “review score” of zero and a missing “review score”, I treat this dummy separately. In total, there are 51,510 sessions in this sample. On average, there are 1.14 clicks per session, and about 8% of sessions ended with a hotel booking. The number of alternatives on the product list of a session varies between 5 and 38. This variation does not result from consumers not browsing further but from these queries being for hotels in destinations or on dates where only a few hotels had available rooms. Some destinations may also offer more alternatives, but the data contains only the results displayed on the first page. As Ursu (2018) notes, this imposes little restriction as position effects are identified from differences across positions. Ursu (2018) also discusses two other limitations, that the data does not contain sessions without a click and that sessions with bookings were over-sampled, and how they can be circumvented.
Table 1: Summary Statistics (Random Ranking)
N Mean Median Std. Dev Min. Max.
Hotel-level
   Price (in \$) 1,357,106 171.70 141.04 114.03 10.00 1000.00
   Star rating 1,333,734 003.34 003 000.89 01 05
   Review score 1,354,996 003.81 004.00 000.97 00.00 05.00
   No reviews 1,354,996 000.04 000 000.19 00 01
   Chain 1,357,106 000.62 001.00 000.48 00.00 01.00
   Location score 1,357,106 003.26 003 001.53 00 07
   On promotion 1,357,106 000.24 000 000.43 00 01
Session-level
   Number of items 51,510 026.35 031 008.46 05 0038
   Number of clicks 51,510 001.14 001 000.66 01 0025
   Made booking 51,510 000.08 000 000.27 00 01
   Trip length (in days) 51,510 003.07 002 002.42 01 0040
   Booking window (in days) 51,510 053.67 031 062.49 00 0498
   Number of adults 51,510 002.08 002 000.94 01 09
   Number of children 51,510 000.43 000 000.82 00 09
   Number of rooms 51,510 001.14 001 000.46 01 08
  • Notes: Summary statistics for sessions under random ranking.
Expedia Benefits from Increasing Total Revenues
By offering hotel bookings on its website, Expedia makes revenues through two different models typical for online travel agents. [^8] Expedia also offers other services, such as car rentals or flights, that are not part of the data. First, Expedia offers hotel stays through a merchant model. Under this model, Expedia negotiates with individual hotels the rooms made available through its website and each room’s respective pricing. Second, Expedia also offers hotel stays through an agency model. Under this model, Expedia takes a commission and only passes on reservations and payments to the individual hotels that set prices themselves. With the data, it is not possible to determine under which revenue model a specific hotel is listed. [^9] In 2013 (the sample period), Expedia made 70% of its worldwide revenue through the merchant model and 24% through the agency model. The income from the merchant model primarily is from hotel bookings, whereas the agency model also includes flights and other products. See the annual report https://s27.q4cdn.com/708721433/files/doc_financials/2013/ar/EXPE_2013_Annual_Report.PDF However, under both models, Expedia makes revenues through a price-based commission, i.e., by taking a share of the price the consumer pays when booking the hotel. As a result, Expedia increases its own revenues by increasing total revenues across all hotels on the list. Moreover, if commissions are the same for all hotels, maximizing total revenues also maximizes Expedia’s profits.
Descriptive Evidence of Heterogeneous Position Effects
I estimate the following linear probability model (LPM) to analyze whether position effects are heterogeneous and how they depend on product attributes:
$\mathbb{P}(Y_{ij}=1|z_{ij})=x_{j}^{\prime}\beta_{1}+w_{i}^{\prime}\beta_{2}+pos_{ij}\gamma+pos_{ij}x_{j}^{\prime}\theta+\tau_{d}\ .$ (2)
Each observation is a hotel $j$ in destination $d$ displayed on position $pos_{ij}=1,2,\dots$ in a consumer query $i$. A larger position means that the hotel is shown further down the list such that a negative coefficient for $\gamma$ implies that hotels further down are less likely to be clicked on or booked. Hotel attributes (e.g., price) are gathered in column vector $x_{j}$, whereas query characteristics (e.g., trip length) are gathered in column vector $w_{i}$. Depending on the specification, $y_{ij}$ is a dummy indicating whether $j$ was clicked on or booked in session $i$. $\tau_{d}$ indicates fixed effects on a destination level. Finally, $z_{ij}$ gathers $x_{j}$, $w_{i}$, $pos_{ij}$ and the fixed effects $\tau_{d}$. Throughout, I denote random variables with capital letters.
The main parameters of interest are $\gamma$ and $\theta$. These parameters jointly capture how a hotel’s position affects its probability of being clicked on or booked. Specifically, given some hotel characteristics $x_{j}$, I define the position effect as the marginal effect of the position on the CTR or booking probability, conditional on hotel attributes:
$\frac{\partial\mathbb{P}(Y_{ij}=1|x_{ij})}{\partial pos_{ij}}=\gamma+x_{j}^{\prime}\theta\ .$ (3)
Given this specification, the position effect captures absolute changes, which are the relevant measures for changes to platform revenues as highlighted in Section 3. Because of the randomized ranking, hotel attributes of all hotels in the list are uncorrelated with any of the positions. Hence, these position effects can be estimated consistently.
Table 2: Coefficient Estimates (LPM, Random Ranking)
Clicks Bookings
(1) (2) (3) (4) (5) (6)
Position -0.1864*** -0.1424*** -0.1427*** -0.0116*** -0.0028 -0.0028
(0.0019) (0.0121) (0.0123) (0.0005) (0.0028) (0.0029)
Price -0.0125*** -0.0173*** -$0**$0011*** -0.0017*** -0.0017***
(0.0002) (0.0004) (0.0004) (0.0000) (0.0001) (0.0001)
On promotion 1.1578*** 1.4961*** 1.5318*** 0.1261*** 0.1436*** 0.1463***
(0.0497) (0.1100) (0.1229) (0.0130) (0.0287) (0.0332)
Star rating 1.6211*** 2.1556*** 0.1051*** 0.1227***
(0.0317) (0.0657) (0.0079) (0.0163)
Review score 0.0810** 0.0287 0.0366*** 0.0794***
(0.0351) (0.0813) (0.0082) (0.0193)
No reviews -0.2488 -0.7767** 0.0872** 0.1412*
(0.1622) (0.3839) (0.0345) (0.0825)
Chain 0.2227*** 0.3105*** 0.0248** 0.0641***
(0.0459) (0.0966) (0.0113) (0.0240)
Location score 0.4381*** 0.5602*** 0.0529*** 0.0556***
(0.0168) (0.0321) (0.0036) (0.0071)
Position $\times$ Price 0.0003*** 0.0003*** 0.0000*** 0.0000***
(0.0000) (0.0000) (0.0000) (0.0000)
Position $\times$ Star rating -0.0318*** -$0**$0011 -0.0005
(0.0028) (0.0029) (0.0007) (0.0007)
Position $\times$ Review score 0.0028 0.0044 -0.0025*** -0.0024***
(0.0035) (0.0035) (0.0008) (0.0008)
Position $\times$ No reviews 0.0295* 0.0338** -0.0032 -0.0026
(0.0164) (0.0166) (0.0035) (0.0036)
Position $\times$ Chain -0.0053 -0.0047 -0.0023** -0.0017
(0.0041) (0.0042) (0.0010) (0.0010)
Position $\times$ Location score -0.0071*** -0.0083*** -0.0002 -0.0004
(0.0013) (0.0013) (0.0003) (0.0003)
Position $\times$ On promotion -0.0201*** -0.0209*** -0.0010 -0.0010
(0.0048) (0.0049) (0.0012) (0.0013)
FE Hotel no no yes no no yes
N 1,220,917 1,220,917 1,219,253 1,220,917 1,220,917 1,219,253
R2 (adj.) 0.0149 0.0152 0.0267 0.0026 0.0026 -0.0019
  • Notes: A larger position means being displayed lower on the product list. Coefficient estimates are scaled to represent changes in terms of percentage points. Estimates for query characteristics and interaction of the squared position with other covariates are omitted. Standard errors are shown in parenthesis and are clustered at the query level. Star ratings are adjusted so that the position effect from the first row is for a hotel with 1 star and other characteristics equal to the minimum observed value.
Table 4.3 presents the results, where coefficient estimates are scaled to represent changes in terms of percentage points. [^10] Standard errors are clustered on a consumer level, capturing that search behavior can induce correlation in the error terms. A large draw in one alternative can mean that consumers are less likely to click on or book another alternative, suggesting a potential negative correlation in the error terms. Columns 1 and 4 show the results of a baseline model that does not include the interaction term and, hence, assumes that position effects are homogeneous. Columns 2 and 5 show the main results. Notably, the interaction coefficient of price and position is positive and significant. This implies that cheaper hotels have stronger position effects on average and conditional on other attributes. For other characteristics, such as the star rating or the review score, the sign of the interaction term also opposes the sign of the direct effect of the characteristics. Hence, desirable attributes that lead to more clicks and bookings on average amplify position effects.
Figure 1 highlights that the resulting differences in position effects are sizable: they increase by about 50% (CTR) or 100% (booking probability) when increasing the price from the 10th to the 90th percentile of the price distribution while keeping other attributes fixed at their respective mean.
Figure 1: Cheaper Hotels Have Stronger Position Effects
Refer to caption (a) Refer to caption (b)
Notes: Position effects as defined in (3) and expressed in percentage points. Other characteristics are set to their respective average. The bars indicate 95% confidence intervals calculated from standard errors clustered on the session level.
Figure 1: Cheaper Hotels Have Stronger Position Effects
Figure 2: Hotel Attributes and Position Effects Complement Each Other
Refer to caption (a) Refer to caption (b)
Notes: Position effects as defined in (3) and expressed in percentage points. Values in $x_{j}$ are set to respective percentile, ordered based on the sign of coefficients in columns 2 and 5 in Table 4.3. The bars indicate 95% confidence intervals calculated from standard errors clustered on the session level.
Figure 2: Hotel Attributes and Position Effects Complement Each Other
Figure 2 shows that there are also sizable differences across different percentiles of the distribution of hotel characteristics $x_{j}$ in the data. Whereas the estimated position effect for clicks is around 0.08 percentage points for a hotel at the 10th percentile, its magnitude doubles for a hotel at the median and more than triples for a hotel at the 90th percentile. Similarly, the position effect for bookings increases from close to zero to around 0.015 percentage points across these percentiles.
While the results so far show that desirable characteristics amplify position effects, they do not yet quantify the degree of heterogeneity in the data. Opposing effects from different attributes may offset each other such that the degree of heterogeneity across hotels observed in the data depends on the joint distribution of attributes. Figure 3 accounts for these potentially opposing effects by depicting the distribution of the estimated average position effects across the hotels in the data. It highlights that position effects differ substantially between hotels in the data. For clicks, the estimated position effects range from below -0.25 to above -0.1, indicating that position effects for some hotels are more than two to three times larger than those of other hotels. Similarly, the estimated position effects for bookings range from below -0.015 to above zero, again highlighting substantial heterogeneity. Though the positive position effects suggest that these hotels could have a demand increase when being moved further down the list, I attribute these few cases to the small size and less precision in coefficient estimates for bookings, as also indicated by the large standard errors reported in Table 4.3.
Figure 3: Position Effects Vary across Hotels in the Data
Refer to caption (a) Refer to caption (b)
Notes: Position effects defined in (3) at hotel attributes $x_{j}$ as observed in the data.
Figure 3: Position Effects Vary across Hotels in the Data